Hi Tutors:
I used the substitution method to solve this system of equation:
m-2n=16
4m+n=1
After going through all the steps my answer is n=-7 and m=30
My problem is I did not get the correct answers when I did a check. Please help me.
SAN ANTONIO is not "hot and dusty, it is hot and humid"

Answer by longjonsilver(2297)

m-2n=16
--> m = 16+2n --eqn1

also 4m+n = 1. Now sub eqn1 into this one, to give

4(16+2n) + n = 1
64 + 8n + n = 1
64 + 9n = 1
9n = -63
n = -63/9
n = -7

hence, from eqn1:
m = 16 + 2(-7) --> this will be your error, here. You will have +14.
m = 16 - 14
m = 2

CHECK:
4m+n
4(2) + (-7)
8 - 7
1
which is correct.

Sharmaine - your working was fine apart from a simple error. You understand the process, so when your check showed you had an error, you go back and you double check EVERYTHING, especially minus signs, as that is usually where students make mistakes.

jon.